OpenStudy (anonymous):

Write an equation in standard form of a circle with the given center and radius. Center (8,1); Radius=5 Will give medals. (show Work please)

OpenStudy (ranga):

The general equation of a circle of radius r centered at (h,k) is: (x - h)^2 + (y - k)^2 = r^2 Just plug in the numbers for h, k and r.

OpenStudy (anonymous):

so, (x-8)^2+(y-1)^2=5^2?

OpenStudy (hoblos):

yes it's correct